if 2(acosx-cos2x)=1 for x belongs to R then find  all the possible


real values of a. 

Grade:

1 Answers

Sunil Raikwar
9 years ago
2acosx-2cos2x-1=0
2acosx-2(2cos2x-1)-1=0
4cos2x+2acosx+1=0
for real roots D greater then equal to zero
4a2-16>=0
a2>=4
so a greater then equal to 2 or less then equal to -2.

Thanks & Regards
Sunil Raikwar
askIITians Faculty